OpenStudy (anonymous):

*Advance Algebra With Financial Application * [WORK IS SHOWN] ***WHAT GET'S ME IS THAT IM STUCK . I SUBTRACTED THE PURCHASE AND THE DOWN PAYMENT WHICH GOT ME :16740 Fitzgerald purchased a car for $18,980. He made a down payment of $2,240. He applied for a five-year installment loan with an interest rate of 7.4%. What is the total cost of the car after five years?

OpenStudy (anonymous):

18,980-2240=16740 M=16740(1+(0.074)/(12))^(12)/(1+(0.074)/(12))^(12)[-1] ?

OpenStudy (amistre64):

installment loan? what are the directives for that?

OpenStudy (amistre64):

is it monthly payments? or just a wait 5 years payment?

OpenStudy (amistre64):

if payments can be calculated, id suggest \[0=B_o(k)^5-P\frac{1-k^5}{1-k}\] \[B_o(k)^5=P\frac{1-k^5}{1-k}\] \[B_o(k)^5\frac{1-k}{1-k^5}=P\]

OpenStudy (amistre64):

where k = (1+r/c)

OpenStudy (amistre64):

and i spose that 5 needs to be adjusted by c as well

OpenStudy (amistre64):

if its monthly payments, that gives me c=12 and, payments of about 4015.68 a month

OpenStudy (amistre64):

\[B_o(1+\frac{.074}{12})^{12*5}\frac{-\frac{.075}{12}}{1-(1+\frac{.074}{12})^{12*5}}=P\]

OpenStudy (anonymous):

O.o where do i start to solve the formula

OpenStudy (amistre64):

lol, im assuming this is just another name for a monthly payment loan that is paid off in 5 years

OpenStudy (amistre64):

Bo = 16740, the rest of it is just how i shuffle the weight around

OpenStudy (amistre64):

lets say k = 1.00617

OpenStudy (amistre64):

\[B_o(1.00617)^{60}~\frac{1-1.00617}{1-1.00617^{60}}=P\]

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=16740%281.00617%29%5E60+*+-.00617%2F%281-1.00617%5E60%29 so payments each month are about 334.68

OpenStudy (amistre64):

if we want to be more exact, 334.65 so not much difference for an approximation

OpenStudy (amistre64):

so about 20079 plus the down payment

OpenStudy (anonymous):

Is that answer ?

OpenStudy (anonymous):

$20,078.40 $22,318.40 $17,978.76 $20,384.52

OpenStudy (amistre64):

it looks like 20078.40 is the payment option if we add back in the 2240 down payment ---------- 22318.40

OpenStudy (anonymous):

@amistre64 Thank youu !!

OpenStudy (amistre64):

youre welcome